OpenStudy (anonymous):

What is the magnitude of the angular acceleration of a 25.0 kg disk of radius .800 m when a net torque is applied to it of 40.0 N-m?

OpenStudy (anonymous):

ok i got it wrong....i got 2.5, but the right answer is 5?? I=25(.8)^2=16 then plug in I so 40/16=angular acc, what did I do wrong?

OpenStudy (shamim):

If a small object rotates around a circle of radius r, then the momentvof inertia I=mr^2

OpenStudy (shamim):

* moment of inertia

OpenStudy (shamim):

This formula will b changed for a rotating disc

OpenStudy (shamim):

I think u hv to use I =0.5mr^2

OpenStudy (shamim):

Response plz!!!!!!!

OpenStudy (anonymous):

where did you get .5?

OpenStudy (anonymous):

oh nevermind! from the formula for a disk, thats what i am missing! thanks a bunch!!!! :D

OpenStudy (shamim):

U r most welcome!!!!!!